# Huppert’s conjecture for finite simple exceptional groups of Lie type

### v1

@@ 5. Verifying Step 2: $G^{\prime}/M$ is isomorphic to $H$
Let $M\leq G^{\prime}$ be a normal subgroup of $G$ such that $G^{\prime}/M$ is a chief factor of $G.$ As $G^{\prime}$ is perfect, $G^{\prime}/M$ is nonabelian so $G^{\prime}/M\cong S^{k}$ for some finite nonabelian simple group $S$ and some integer $k\geq 1.$
5.1. Eliminating the alternating groups of degree at least $7$
Assume that $S=\mathrm{A}_{n},n\geq 7.$ Let $\theta_{i},i=1,2$ be irreducible characters of $S$ obtained from Lemma 2.1$(i)$ with $\theta_{1}(1)=n(n-3)/2$ and $\theta_{2}(1)=(n-1)(n-2)/2.$ We have that both $\theta_{i}$ extend to ${\mathop{\rm Aut}\nolimits}(\mathrm{A}_{n})\cong\mathrm{S}_{n}$ and $\theta_{2}(1)=\theta_{1}(1)+1.$ By [2, Lemma 5], we obtain $\theta_{i}^{k}\in{\mathrm{Irr}}(G^{\prime}/M)$ extends to $G/M,$ for $i=1,2.$ Hence $\theta_{i}(1)^{k}\in{\mathrm{cd}}(G)$ and $\theta_{i}(1)^{k}$ are coprime for $i=1,2.$ Note that $\mathrm{S}_{n}$ also has an irreducible character labelled by the partition $(n-1,1)$ which is irreducible upon restriction to $\mathrm{A}_{n}$. Thus $\mathrm{A}_{n}$ has an irreducible character $\theta$ of degree $n-1$ which is extendible to $\mathop{\rm Aut}\nolimits(\mathrm{A}_{n})\cong\mathrm{S}_{n}$. Hence $\theta(1)^{k}=(n-1)^{k}\in{\mathrm{cd}}(G)$.
We will need the following observation:
@@ 5.1. Eliminating the alternating groups of degree at least $7$
Subcase $H\cong\textrm{F}_{4}(q),q>2.$ Assume first that $p\neq 3.$ By Lemma 3.1$(viii),$ one of the degrees $\theta_{i}(1)^{k},i=1,2,$ must be $q^{24}.$ However we have that $(n-1,n-2)=1$ and $(n,n-3)=(n,3)$ so $\theta_{i}(1)^{k}$ can never be a power of $p\neq 3.$ Thus $p=3.$ If $\theta_{j}^{k}(1)=q^{24}$ for some $j,$ then $j=1$ and hence $\theta_{1}^{k}(1)=(n(n-3)/2)^{k}=3^{24f},$ where $q=3^{f}.$ As $n\geq 7,$ we see that $n(n-3)/2$ is a prime power if and only if $n=9,$ so that $\theta_{1}^{k}(1)=3^{3k}=3^{24f}.$ Now $G$ has a degree $(n-1)^{k}=8^{k}$. By [18, Theorem $1.1$], the only nontrivial prime power degree of $G$ is $q^{24}=3^{24f},$ whence $3^{24f}=8^{k}$, a contradiction. Now Lemma 3.1$(viii)$ yields
@@
By Lemma 3.1$(ix)$, $G$ has no consecutive degrees so $k\geq 2.$ But then $\Phi_{3}\Phi_{6}\Phi_{12}$ is a proper nontrivial power, which is impossible by Lemma 2.2.
@@
Subcase $H\cong{}^{2}\textrm{E}_{6}(q).$ Assume first that $p\neq 3.$ By Lemma 3.2$(viii),$ one of the degrees $\theta_{i}(1)^{k},i=1,2,$ must be $q^{36}.$ However we have that $(n-1,n-2)=1$ and $(n,n-3)=(n,3)$ so $\theta_{i}(1)^{k}$ can never be a power of $p\neq 3.$ Thus $p=3.$ If $\theta_{j}^{k}(1)=q^{36}$ for some $j,$ then $j=1$ and hence $\theta_{1}^{k}(1)=(n(n-3)/2)^{k}=3^{36f},$ where $q=3^{f}.$ As $n\geq 7,$ we see that $n(n-3)/2$ is a prime power if and only if $n=9,$ so that $\theta_{1}^{k}(1)=3^{3k}=3^{36f}.$ Again, $G$ has a degree $(n-1)^{k}=8^{k}$. However, by [18, Theorem $1.1$], the only nontrivial prime power degree of $G$ is $q^{36}=3^{36f},$ we obtain a contradiction. Thus by Lemma 3.2$(viii)$, we have $\{\theta_{1}^{k}(1),\theta_{2}^{k}(1)\}=\{\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18},\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}\}.$ By Lemma 3.2$(ix)$, $G$ has no consecutive degrees so $k\geq 2.$ But then $\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18}$ is a proper nontrivial power. As $q=3^{f},$ we deduce that $\Phi_{3}$ is coprime to $\Phi_{6}^{2}\Phi_{12}\Phi_{18}$ and so $\Phi_{3}$ must be a nontrivial proper power, which is impossible by Lemma 2.2.
The arguments for the remaining groups $\textrm{E}_{6}(q),\textrm{E}_{7}(q$ and $\textrm{E}_{8}(q)$ are similar.
5.2. Eliminating the sporadic simple groups and the Tits group.
@@
Suppose that $S$ is a sporadic simple group or the Tits group. By Table 1, $S$ possesses two nontrivial irreducible characters $\theta_{i},i=1,2,$ such that $\theta_{i}(1)^{\prime}s$ are coprime and each $\theta_{i}$ extends to $\mathop{\rm Aut}\nolimits(S).$ By [2, Lemma 5], for each $i,$ $\theta_{i}^{k}$ extends to $G,$ so $\theta_{i}^{k}(1)\in{\mathrm{cd}}(G).$ Observe that for each $i,$ $\theta_{i}(1)$ is not prime power so $\theta_{i}^{k}(1)\neq|H|_{p}.$
Assume $H\cong{\rm F}_{4}(q)$. As $(\theta_{1}^{k}(1),\theta_{2}^{k}(1))=1,$ we deduce from Lemma 3.1$(viii)$ that $q=3^{f}$ and $\{\theta_{1}^{k}(1),\theta_{2}^{k}(1)\}=\{\Phi_{3}\Phi_{6}\Phi_{12},\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}\}.$ In this situation, the larger degree among $\theta_{i}(1),i=1,2,$ must be divisible by $2^{5}\cdot 3\cdot 5\cdot 41$ However this is not the case by inspecting the degrees in Table 1.
@@
Assume $H\cong{}^{2}{\rm E}_{6}(q)$. As $(\theta_{1}^{k}(1),\theta_{2}^{k}(1))=1,$ we deduce from Lemma 3.2$(viii)$ that $q=3^{f}$ and $\{\theta_{1}^{k}(1),\theta_{2}^{k}(1)\}=\{\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18},\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}\}.$ In this situation, one of the degrees $\theta_{i}(1),i=1,2,$ must be divisible by $2^{5}\cdot 3\cdot 5\cdot 41$ However this is not the case by inspecting the degrees in Table 1.
@@
Assume $H\cong{\rm E}_{6}(q)$. As $(\theta_{1}^{k}(1),\theta_{2}^{k}(1))=1,$ we deduce from Lemma 3.3$(viii)$ that $q=3^{f}$ and $\{\theta_{1}^{k}(1),\theta_{2}^{k}(1)\}=\{\Phi_{3}^{2}\Phi_{6}\Phi_{9}\Phi_{12},\frac{1}{3}q^{7}\Phi_{1}^{6}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{5}\Phi_{8}\}.$ In this situation, one of the degrees $\theta_{i}(1),i=1,2,$ must be divisible by $2^{5}\cdot 3\cdot 5\cdot 41$ However this is not the case by inspecting the degrees in Table 1.
@@
Assume $H\cong{\rm E}_{7}(q)$ or ${\rm E}_{8}(q)$. Observe that both $\theta_{i}(1)$ are not prime powers so both $\theta_{i}^{k}(1)$ are pairwise coprime and different from $|H|_{p},$ which contradicts Lemma 3.4$(viii)$ and Lemma 3.5$(viii)$, respectively.
5.3. Reduction to finite simple groups of Lie type in characteristic $p$
@@
In this subsection, we show that if $S$ is a finite simple group of Lie type in characteristic $r,$ with $S\neq{}^{2}{\rm F}_{4}(2)^{\prime},$ then $k=1$ and $r=p.$ Let $\theta$ be the Steinberg character of $S.$ Then $\theta(1)=|S|_{r}$ and $\theta$ extends to ${\mathop{\rm Aut}\nolimits}(S).$ By [2, Lemma 5], $\theta^{k}\in{\mathrm{Irr}}(G^{\prime}/M)$ extends to $G/M,$ hence $\theta(1)^{k}=|S|_{r}^{k}\in{\mathrm{cd}}(G).$ Since $|H|_{p}$ is the unique nontrivial prime power character degree of $G$ by [18, Theorem $1.1$], we have $\theta(1)^{k}=|H|_{p}.$ In particular, we have $r=p.$
@@
We next claim that $k=1$. By way of contradiction, assume that $k\geq 2.$ Write $\theta(1)=q_{1}^{s}$ and $|H|_{p}=q^{a(H)}$, so $a(H)=24,36,36,63,120$, respectively (see [15, Table 5.1.B]). Then $q_{1}^{sk}=q^{a(H)}.$ Let $\psi=\tau\times\theta\times\cdots\times\theta\in{\mathrm{Irr}}(S^{k}),$ where $\tau\in{\mathrm{Irr}}(S)$ with $1<\tau(1)\neq\theta(1).$ Then $\psi(1)=\tau(1)q_{1}^{s(k-1)}\nmid q^{a(H)}$ and is nontrivial, so it must divide some character degree of $G,$ which is different from $q^{a(H)}.$ By part $(vi)$ of Lemmas 3.13.5, we have that $q_{1}^{s(k-1)}=q^{a(H)(k-1)/k}\leq q^{b(H)}$, where $b(H)=16,25,25,46,91$, respectively, and hence $a(H)(k-1)\leq kb(H),$ which implies that $k\leq k_{0}$ for some positive integer $k_{0}$. We can check that $k_{0}=3$ for all groups except for ${\rm E}_{8}(q)$, where $k_{0}=4$.
@@
Let $C$ be a normal subgroup of $G$ such that $C/M=C_{G/M}(G^{\prime}/M).$ Then $G^{\prime}C/C\cong S^{k}$ is a unique minimal normal subgroup of $G/C$ so $G/C$ embeds into ${\mathop{\rm Aut}\nolimits}(S)\wr S_{k},$ where $S_{k}$ is the symmetric group of degree $k.$ Let $B={\mathop{\rm Aut}\nolimits}(S)^{k}\cap G/C.$ Then $|G/C:B|\mid k!.$ Let $\psi=1\times\theta\times\cdots\times\theta\in{\mathrm{Irr}}(G^{\prime}C/C).$ Then $\psi$ extends to $\psi_{0}\in{\mathrm{Irr}}(B)$ by using the representation theory of wreath products. Let $J=I_{G}(\psi_{0})$ and let $\mu\in{\mathrm{Irr}}(J|\psi_{0}).$ By [13, Theorem 6.11], $\chi=\mu^{G}\in{\mathrm{Irr}}(G).$ Then $\chi(1)=e\psi_{0}(1)=e\psi(1)=eq^{a(H)(k-1)/k},$ where $e\mid k!.$ Since $k\leq 4,$ we deduce that $e\mid 4!=24,$ so $1<\chi(1)<q^{a(H)}.$ In addition $\chi(1)$ is divisible by $q^{\lfloor a(H)/2\rfloor}.$ As $\ell_{1}$ and $\ell_{2}$ are primitive prime divisors of $\Phi_{m_{1}}$ and $\Phi_{m_{2}},$ where $m_{1}>m_{2}$ are the two largest integers such that $\Phi_{m_{i}}$ divides $|H|$, it follows from Lemma 2.5 that $\ell_{1}\geq m_{1}+1$ and $\ell_{2}\geq m_{2}+1,$ in particular, $\ell_{1},\ell_{2}\geq 11$. So $(\ell_{1}\ell_{2},\chi(1))=1,$ and then by part $(i)$ of Lemmas 3.13.5, we obtain that $\chi(1)=\frac{1}{t}q^{c(H)}\alpha,$ where $1\leq t\leq 6$ is an integer, $c(H)=4,7,7,11,16$, respectively and $\alpha$ is an integer coprime to $p$. However, this is a contradiction since $\chi(1)$ is divisible by $q^{\lfloor a(H)/2\rfloor}>q^{c(H)}$. Thus $k=1$ and $r=p.$
5.4. Eliminating finite simple groups of Lie type in characteristic $p$.
@@
We will show that if $G^{\prime}/M\cong S$ is a simple group of Lie type in characteristic $p$ and $S\neq{}^{2}{\rm F}_{4}(2)^{\prime},$ then $S\cong H.$ We will prove this by eliminating other possibilities of $S.$ Assume that $S$ is a simple group of Lie type in characteristic $p$ and $S$ is not the Tits group. We have shown that $G^{\prime}/M\cong S$ and $|S|_{p}=|H|_{p}=q^{a(H)}.$ Observe that if $\theta\in{\mathrm{Irr}}(S)$ is extendible to ${\mathop{\rm Aut}\nolimits}(S),$ then $\theta$ extends to $G/C,$ where $C/M=C_{G/M}(G^{\prime}/M),$ so $\theta(1)\in{\mathrm{cd}}(G/C)\subseteq{\mathrm{cd}}(G).$ In fact, we will choose $\theta$ to be a unipotent character of $S,$ so by results of Lusztig, $\theta$ is extendible to ${\mathop{\rm Aut}\nolimits}(S)$ apart from some exceptions (see [17, Theorem $2.5$]). We refer to [4, $13.8,13.9$] for the classification of unipotent characters and the notion of symbols. In Table 2, for each simple group of Lie type $S$ in characteristic $p,$ we list the $p$-part of some unipotent character of $S$ that is extendible to ${\mathop{\rm Aut}\nolimits}(S).$

### v3

@@ 5. Verifying Step 2: $G^{\prime}/M$ is isomorphic to $H$
Let $H$ be one of the simple groups listed in (1) and let $G$ be a finite group such that ${\mathrm{cd}}(G)={\mathrm{cd}}(H)$. By Proposition 4.1, $G^{\prime}=G^{\prime\prime}$. Let $M\leq G^{\prime}$ be a normal subgroup of $G$ such that $G^{\prime}/M$ is a chief factor of $G$. The main purpose of this section is to show that $G^{\prime}/M\cong H.$ Recall that $H$ is a finite simple exceptional group of Lie type defined over a field of size $q=p^{f}$, where $p$ is a prime.
Proposition 5.1.
Let $G$ be a finite group and let $H$ be one of the finite simple groups in (1). Let $M\leq G^{\prime}$ be a normal subgroup of $G$ such that $G^{\prime}/M$ is a chief factor of $G$. Then $G^{\prime}/M\cong H.$
Proof.
By Proposition 4.1, $G^{\prime}$ is perfect. Hence $G^{\prime}/M\cong S^{k}$, where $S$ is a nonabelian simple group and $k\geq 1$ is an integer. Using the classification of finite simple groups, we first show that $S$ must be a finite simple groups of Lie type in characteristic $r$ for some prime $r$ and then show that $r=p,k=1$ and $S\cong H$.
(1) Assume $S=\mathrm{A}_{n},n\geq 7.$ Let $\theta_{i},i=1,2,$ be irreducible characters of $S$ obtained from Lemma 2.1$(i)$ with $\theta_{1}(1)=n(n-3)/2$ and $\theta_{2}(1)=(n-1)(n-2)/2.$ Now both $\theta_{i}$ extend to ${\mathop{\rm Aut}\nolimits}(\mathrm{A}_{n})\cong\mathrm{S}_{n}$ and $\theta_{2}(1)=\theta_{1}(1)+1.$ By Lemma 2.2, $\theta_{i}^{k}\in{\mathrm{Irr}}(G^{\prime}/M)$ extends to $G/M,$ for $i=1,2.$ Hence $\theta_{i}(1)^{k}\in{\mathrm{cd}}(G)$ and $\theta_{i}(1)^{k}$ are coprime for $i=1,2.$ Note that $\mathrm{S}_{n}$ also has an irreducible character labelled by the partition $(n-1,1)$ which is irreducible upon restriction to $\mathrm{A}_{n}$. Thus $\mathrm{A}_{n}$ has an irreducible character $\theta$ of degree $n-1$ which is extendible to $\mathop{\rm Aut}\nolimits(\mathrm{A}_{n})\cong\mathrm{S}_{n}$. Hence $\theta(1)^{k}=(n-1)^{k}\in{\mathrm{cd}}(G)$.
We need the following observation:
@@
Subcase $H\cong\textrm{F}_{4}(q),q>2.$ Assume first that $p\neq 3.$ By Lemma 3.2$(viii),$ one of the degrees $\theta_{i}(1)^{k},i=1,2,$ must be $q^{24}.$ However, $(n-1,n-2)=1$ and $(n,n-3)=(n,3)$ so $\theta_{i}(1)^{k}$ can never be a power of $p\neq 3.$ Thus $p=3.$ If $\theta_{j}(1)^{k}=q^{24}$ for some $j,$ then $j=1$ and hence $\theta_{1}(1)^{k}=(n(n-3)/2)^{k}=3^{24f},$ where $q=3^{f}.$ As $n\geq 7,$ $n(n-3)/2$ is a prime power if and only if $n=9,$ so that $\theta_{1}(1)^{k}=3^{3k}=3^{24f}.$ Now $G$ has a degree $(n-1)^{k}=8^{k}$. By Lemma 2.4, the only nontrivial prime power degree of $G$ is $|H|_{p}=q^{24}=3^{24f},$ whence $3^{24f}=8^{k}$, a contradiction. Lemma 3.2$(viii)$ yields
@@
By Lemma 3.2$(ix)$, $G$ has no consecutive degrees, so $k\geq 2.$ But then $\Phi_{3}\Phi_{6}\Phi_{12}$ is a proper nontrivial power, which is impossible by Lemma 2.3.
@@
Subcase $H\cong{}^{2}\textrm{E}_{6}(q).$ Assume first that $p\neq 3.$ By Lemma 3.1$(viii),$ one of the degrees $\theta_{i}(1)^{k},i=1,2,$ must be $q^{36}.$ However $(n-1,n-2)=1$ and $(n,n-3)=(n,3)$ so $\theta_{i}(1)^{k}$ can never be a power of $p\neq 3.$ Thus $p=3.$ If $\theta_{j}(1)^{k}=q^{36}$ for some $j,$ then $j=1$ and hence $\theta_{1}(1)^{k}=(n(n-3)/2)^{k}=3^{36f},$ where $q=3^{f}.$ As $n\geq 7,$ $n(n-3)/2$ is a prime power if and only if $n=9,$ so that $\theta_{1}(1)^{k}=3^{3k}=3^{36f}.$ Again, $G$ has a degree $(n-1)^{k}=8^{k}$. However, by Lemma 2.4, the only nontrivial prime power degree of $G$ is $q^{36}=3^{36f},$ so we obtain a contradiction. Thus by Lemma 3.1$(viii)$,
$\{\theta_{1}(1)^{k},\theta_{2}(1)^{k}\}=\{\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18},\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}\}.$
By Lemma 3.1$(ix)$, $G$ has no consecutive degrees so $k\geq 2.$ But then $\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18}$ is a proper nontrivial power. As $q=3^{f},$ $\Phi_{3}$ is coprime to $\Phi_{6}^{2}\Phi_{12}\Phi_{18}$ and so $\Phi_{3}$ is a nontrivial proper power, which is impossible by Lemma 2.3.
The arguments for the remaining groups $\textrm{E}_{6}(q),\textrm{E}_{7}(q)$ and $\textrm{E}_{8}(q)$ are similar.
@@
(2) Suppose that $S$ is a sporadic simple group or the Tits group. By Table 1, $S$ possesses two nontrivial irreducible characters $\theta_{i},i=1,2,$ such that $\theta_{1}(1)$ and $\theta_{2}(1)$ are coprime and each $\theta_{i}$ extends to $\mathop{\rm Aut}\nolimits(S).$ By Lemma 2.2, for each $i,$ $\theta_{i}^{k}$ extends to $G,$ so $\theta_{i}(1)^{k}\in{\mathrm{cd}}(G).$ Observe that for each $i,$ $\theta_{i}(1)$ is not a prime power so $\theta_{i}(1)^{k}\neq|H|_{p}.$
Assume $H\cong{\rm F}_{4}(q)$. As $(\theta_{1}(1)^{k},\theta_{2}(1)^{k})=1,$ by Lemma 3.2$(viii)$, $q=3^{f}$ and
$\{\theta_{1}(1)^{k},\theta_{2}(1)^{k}\}=\{\Phi_{3}\Phi_{6}\Phi_{12},\frac{1}{3}q^{4}\Phi_{1}^{4}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{8}\}.$
@@
In this situation, the larger of $\theta_{i}(1)$ and $\theta_{2}(1)$ must be divisible by $2^{5}\cdot 3\cdot 5\cdot 41$ However this is not the case by inspecting the degrees in Table 1.
Assume $H\cong{}^{2}{\rm E}_{6}(q)$. As $(\theta_{1}(1)^{k},\theta_{2}(1)^{k})=1,$ Lemma 3.1$(viii)$ yields that $q=3^{f}$ and
$\{\theta_{1}(1)^{k},\theta_{2}(1)^{k}\}=\{\Phi_{3}\Phi_{6}^{2}\Phi_{12}\Phi_{18},\frac{1}{3}q^{7}\Phi_{1}^{4}\Phi_{2}^{6}\Phi_{4}^{2}\Phi_{8}\Phi_{10}\}.$
@@
In this situation, one of the degrees $\theta_{i}(1),i=1,2,$ must be divisible by $2^{5}\cdot 3\cdot 5\cdot 41$ However this is not the case by inspecting the degrees in Table 1.
Assume $H\cong{\rm E}_{6}(q)$. As $(\theta_{1}(1)^{k},\theta_{2}(1)^{k})=1,$ Lemma 3.3$(viii)$ implies that $q=3^{f}$ and
$\{\theta_{1}(1)^{k},\theta_{2}(1)^{k}\}=\{\Phi_{3}^{2}\Phi_{6}\Phi_{9}\Phi_{12},\frac{1}{3}q^{7}\Phi_{1}^{6}\Phi_{2}^{4}\Phi_{4}^{2}\Phi_{5}\Phi_{8}\}.$
In this situation, one of the degrees $\theta_{i}(1),i=1,2,$ must be divisible by $2^{5}\cdot 3\cdot 5\cdot 41$ However this is not the case by inspecting the degrees in Table 1.
@@
Assume $H\cong{\rm E}_{7}(q)$ or ${\rm E}_{8}(q)$. Observe that both $\theta_{i}(1)$ are not prime powers so both $\theta_{1}(1)^{k}$ and $\theta_{2}(1)^{k}$ are pairwise coprime and different from $|H|_{p},$ which contradicts Lemma 3.4$(viii)$ and Lemma 3.5$(viii)$, respectively.
@@
(3) Assume that $S$ is a finite simple group of Lie type in characteristic $r,$ with $S\neq{}^{2}{\rm F}_{4}(2)^{\prime}.$ Let $\theta$ be the Steinberg character of $S.$ Then $\theta(1)=|S|_{r}$ and $\theta$ extends to ${\mathop{\rm Aut}\nolimits}(S).$ By Lemma 2.2, $\theta^{k}\in{\mathrm{Irr}}(G^{\prime}/M)$ extends to $G/M,$ hence $\theta(1)^{k}=|S|_{r}^{k}\in{\mathrm{cd}}(G).$ Since $|H|_{p}$ is the unique nontrivial prime power character degree of $G$ by Lemma 2.4, $\theta(1)^{k}=|H|_{p}.$ In particular, $r=p.$
@@
We next claim that $k=1$. By way of contradiction, assume $k\geq 2.$ Write $\theta(1)=q_{1}^{s}$ and $|H|_{p}=q^{a(H)}$, so $a(H)=24,36,36,63,120$, respectively (see [18, Table 5.1.B]). Then $q_{1}^{sk}=q^{a(H)}.$ Let $\psi=\tau\times\theta\times\cdots\times\theta\in{\mathrm{Irr}}(S^{k}),$ where $\tau\in{\mathrm{Irr}}(S)$ with $1<\tau(1)\neq\theta(1).$ Then $\psi(1)=\tau(1)q_{1}^{s(k-1)}\nmid q^{a(H)}$ and is nontrivial, so it divides some character degree of $G,$ which is different from $q^{a(H)}.$ By part $(vi)$ of Lemmas 3.13.5, $q_{1}^{s(k-1)}=q^{a(H)(k-1)/k}\leq q^{b(H)}$, where $b(H)=25,16,25,46,91$, respectively, and hence $a(H)(k-1)\leq kb(H),$ which implies that $k\leq k_{0}$ for some positive integer $k_{0}$. We check that $k\leq k_{0}$, where $k_{0}=3$ unless $H={\rm E}_{8}(q)$ where $k_{0}=4$.
@@
Let $C$ be a normal subgroup of $G$ such that $C/M=C_{G/M}(G^{\prime}/M).$ Then $G^{\prime}C/C\cong S^{k}$ is a unique minimal normal subgroup of $G/C$, so $G/C$ embeds into ${\mathop{\rm Aut}\nolimits}(S)\wr S_{k},$ where $S_{k}$ is the symmetric group of degree $k.$ Let $B={\mathop{\rm Aut}\nolimits}(S)^{k}\cap G/C.$ Then $|G/C:B|\mid k!.$ Let $\psi=1\times\theta\times\cdots\times\theta\in{\mathrm{Irr}}(G^{\prime}C/C).$ Then $\psi$ extends to $\psi_{0}\in{\mathrm{Irr}}(B)$. Let $J=I_{G}(\psi_{0})$ and let $\mu\in{\mathrm{Irr}}(J|\psi_{0}).$ By [16, Theorem 6.11], $\chi=\mu^{G}\in{\mathrm{Irr}}(G).$ Then
$\chi(1)=e\psi_{0}(1)=e\psi(1)=eq^{a(H)(k-1)/k},$
where $e\mid k!.$ Since $k\leq 4,$ we deduce that $e\mid 4!=24,$ so $1<\chi(1)<q^{a(H)}.$ In addition $\chi(1)$ is divisible by $q^{\lfloor a(H)/2\rfloor}.$ As $\ell_{1}$ and $\ell_{2}$ are primitive prime divisors of $\Phi_{m_{1}}$ and $\Phi_{m_{2}},$ where $m_{1}>m_{2}$ are the two largest integers such that $\Phi_{m_{i}}$ divides $|H|$, it follows from Lemma 2.7 that $\ell_{1}\geq m_{1}+1$ and $\ell_{2}\geq m_{2}+1$. In particular, $\ell_{1},\ell_{2}\geq 11$. So $(\ell_{1}\ell_{2},\chi(1))=1.$ By part $(i)$ of Lemmas 3.13.5, $\chi(1)=\frac{1}{t}q^{c(H)}\alpha,$ where $t=3,4,3,2,6$ and $c(H)=7,4,7,11,16$, respectively and $\alpha$ is an integer coprime to $p$. However, this is a contradiction since $\chi(1)$ is divisible by $q^{\lfloor a(H)/2\rfloor}>q^{c(H)}$. Thus $k=1$ and $r=p.$
@@
(4) Eliminating finite simple groups of Lie type in characteristic $p$. We show that if $G^{\prime}/M\cong S$ is a simple group of Lie type in characteristic $p$ and $S\neq{}^{2}{\rm F}_{4}(2)^{\prime},$ then $S\cong H.$ We prove this by eliminating other possibilities of $S.$ Assume that $S$ is a simple group of Lie type in characteristic $p$ and $S$ is not the Tits group. We have shown that $G^{\prime}/M\cong S$ and $|S|_{p}=|H|_{p}=q^{a(H)}.$ Observe that if $\theta\in{\mathrm{Irr}}(S)$ is extendible to ${\mathop{\rm Aut}\nolimits}(S),$ then $\theta$ extends to $G/C,$ where $C/M=C_{G/M}(G^{\prime}/M),$ so $\theta(1)\in{\mathrm{cd}}(G/C)\subseteq{\mathrm{cd}}(G).$ In fact, we choose $\theta$ to be a unipotent character of $S.$ By [21, Theorem $2.5$], $\theta$ is extendible to ${\mathop{\rm Aut}\nolimits}(S)$ apart from some exceptions. We refer to [4, $13.8,13.9$] for the classification of unipotent characters and the notion of symbols. In Table 2, for each simple group of Lie type $S$ in characteristic $p,$ we list the $p$-part of some unipotent character of $S$ that is extendible to ${\mathop{\rm Aut}\nolimits}(S).$